OpenStudy (anonymous):

The difference between three times a number and seven is twenty. What is the number?

OpenStudy (vijeya3):

Let the unknown number be x now 3 times that number wud be... 3x next,we gotta subtract 7 from it... so it would be...(3x-7) It is given the value of this number is 20 therefore... (3x-7)=20 3x=20+7 3x=27 x=9 I hope this is helpful.. If it is,fan and medal me.. and tag me whenever you need help:)